Compute the minimum passing sight distance for the following data:
Speed of the passing car = 90 kph
Speed of the overtaken vehicle = 80 kph
Time of initial maneuver = 4 sec
Time passing vehicle occupies the left lane = 9 sec
Distance between the passing vehicle at the end of its maneuver and the opposing vehicle = 80 m
1
Expert's answer
2021-10-15T12:42:14-0400

Passing sight distance

"D=d_1+d_2+d_3\\
d_1=0.278 v_b*T_1=0.278*90*4=88.08 m\\
d_2=0.278 v_b*T_2+ \frac{1}{2}*a*T_2=0.278*90*4+ \frac{1}{2}*2.5*9=211.41 m\\
d_3=80 m\\
D=d_1+d_2+d_3=88.08+211.41+80=379.49 m\\"